OpenStudy (abbie):

Determine the equation of the line that passes through the points (2, -5) and (7, 0). Use the form y=mx+b.

OpenStudy (abbie):

A. y=5x+5 B. y=x-7 C. y=2x-5 D. y=x+7

OpenStudy (kamibug):

Do you want me to show you how to solve it (recommended) or to give you a straight answer?

OpenStudy (anonymous):

Cant give answers straight out you can get beaned

OpenStudy (kamibug):

Okay. I don;t know what "beaned" means but it doesn't sound pleasant. LOL OKay, so let me show you the steps. It's not as hard as you think.

OpenStudy (abbie):

banned he means""

OpenStudy (anonymous):

sorry XD typo

OpenStudy (abbie):

but reaper u dont have to be that harsh

OpenStudy (anonymous):

No just letting u know i got banned before by doin that

OpenStudy (kamibug):

Haha lol, okay so the first thing you need to know is the slope. YOu can find the slope using the points given and the slope formula. http://0.tqn.com/d/create/1/0/9/p/C/-/slopeformula.jpg

OpenStudy (abbie):

ok

OpenStudy (anonymous):

Familiar with the equation 

OpenStudy (kamibug):

Let me know when you have the slope

OpenStudy (anonymous):

kami is faster than me xD

OpenStudy (anonymous):

http://www.webmath.com/equline1.html i use to use this when i didnt understand this

OpenStudy (kamibug):

cool site

OpenStudy (anonymous):

Yeah haha

OpenStudy (kamibug):

thnx for the medal reaper. ill keep in mind the thing about getting banned and giving straight answers lol i din't know. butya gotta listen to those with experience lol

OpenStudy (anonymous):

Np ...i be back need to help my lil sis in her school